OpenStudy (anonymous):

A while back, either James borrowed $12 from his friend Rita or she borrowed $12 from him, but he can’t quite remember which. Either way, he is planning to pay her back or ask that she pay him back this afternoon. If he has $42.80 in his wallet, which equation represents the amounts of money he may have after he sees Rita? a. |x – 42.80| = 12 b. |x + 42.80| = 12 c. |x – 12| = 42.80 d. |x + 12| = 42.80

OpenStudy (anonymous):

@shortycme

OpenStudy (anonymous):

either A or C

OpenStudy (anonymous):

@nirmalnema is one of those correct ?

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@gorv

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\text{u have to choose option here}\\~\\ \text{such that x=42.80+12=54.80 }\\~\\ \text{or x=42.80-12=30.80}}\end{align}\)

OpenStudy (anonymous):

i don't understand? Can you explain it a little more please :)

OpenStudy (mathmath333):

either James borrowed $12 from his friend Rita or she borrowed $12 from him so either james will gain $ 12 or he will give $12

OpenStudy (anonymous):

So would it be C ?

OpenStudy (mathmath333):

No

OpenStudy (anonymous):

So James borrowed $12 from rita ?

OpenStudy (anonymous):

dont feel bad justsmile im confused to

OpenStudy (mathmath333):

for First option lets check \(\large \begin{align} \color{black}{ |x – 42.80| = 12\\~\\ \pm(x – 42.80) = 12\\~\\ (x – 42.80) = 12 ~~or~~(x – 42.80) = -12\\~\\ x =42.80+ 12 ~~or~~x = 42.80 -12\\~\\ x =42.80+ 12 ~~or~~x = 42.80 -12\\~\\ x =54.80 ~~or~~x = 30.80 }\end{align}\)

OpenStudy (mathmath333):

so option A is the correct one here

OpenStudy (anonymous):

thank you . I understand now :)